OpenStudy (anonymous):

Paraobla Help I have a parabola plotted on a graph, but I do not understand how to put it into an equation or find its vertex. I have the two points it starts and stops on tho. (0,0) and (10,0)

OpenStudy (anonymous):

the parabola can open upwards or downwards... which way does it open?

OpenStudy (anonymous):

it goes up

OpenStudy (anonymous):

what do you mean?

OpenStudy (zepp):

(0,0) and (10,0) These things looks like the x-intercepts

OpenStudy (zepp):

Give us another random point on the parabola, and we can solve it :]

OpenStudy (anonymous):

ohh okay holdon

OpenStudy (anonymous):

Theres a picture of it. (5,2) is where it goes to about

OpenStudy (zepp):

Vertex is approximately at (5.5, 3)

OpenStudy (zepp):

x-intercepts at (1,0) and (10,0)

OpenStudy (zepp):

By using the root form (or whatever this is) factored form I believe; \(\large y=a(x-z_1)(x-z_2)\) Where a is a constant that determine the fatness and direction of the parabola z1 and z2 are x-intercepts (or roots)

OpenStudy (zepp):

Let (x,y) be the vertex; (5.5, 3) Then \(\large y = a(x-z_1)(x-z_2)\) Plug in stuffs \(\large 3 = a(5.5-1)(5.5-10)\) Find a \(\large 3 = a(4.5)(-4.5)\) \(\large a = -\frac{4}{27}\)

OpenStudy (zepp):

The equation of this parabola would be \(\large y=-\frac{4}{27}(x-1)(x-10)\)

OpenStudy (anonymous):

i need to use this formula for the equation think of a quadratic equation in general form, y = a(x - h)2 + k

OpenStudy (anonymous):

what is a? and what are h and k?

OpenStudy (zepp):

(h,k) are the vertex

OpenStudy (zepp):

What's your vertex?

OpenStudy (anonymous):

Vertex is approximately at (5.5, 3)

OpenStudy (zepp):

What's a?

OpenStudy (anonymous):

I dont know

OpenStudy (zepp):

http://puu.sh/yoeQ I'm a little bit angry about the fact that you didn't even take the time to read my explanations :|